2. [Theory] For the following questions, the present is Year 0, and the relevant interest rate is 10% per year. Round your answers to the nearest cent. Show your work, or you will not receive marks. There’s a formula sheet at the end of the assignment.

a. (2 marks) Use (P/F,i,N) to find the present value (Year 0 value) of $1,000 in Year 10.

Present Value: $ _______________

b. (4 marks) Maintenance costs for a machine are $100 in Year 1, after which they go up by $50 a year. Thus, costs are $150 in year 2, $200 in Year 3, etc.  The last maintenance costs are incurred in Year 30.  Use (P/A,i,N) and (A/G,i,N) to find the present value (Year 0 value) of these maintenance costs.

Present Value: $ _______________

c. (4 marks) An after-school sports program for disadvantaged children provides benefits worth $500 in Year 1. Each year after that, the benefits from the program go up by 10%, so they are $550 in Year 2, $605 in Year 3, etc. The last year in which benefits are provided is Year 20. Use (P/A,g,i,N) (or equivalently, (P/A,io,N)/(1+g) ) to calculate the present value of these benefits.

Present Value: $_____________

d. (2 marks) You want to save up $1,000,000 by the time you retire in Year 45.  Use (A/F,i,N) to calculate how much money you must put in a bank account every year, starting in Year 1 and with the last payment being in Year 45, in order to have $1,000,000 in Year 45.

Annual Payment: $______________

(F/P,i,N) = (1+i)N

(P/F,i,N) = (1+i)−N

(A/F,i,N) = i(1+i)N−1

(F/A,i,N) = [(1+i)N−1]i

(A/P,i,N) = i(1+i)N(1+i)N−1

(P/A,i,N) = (1+i)N−1i(1+i)N

(A/G,i,N) = 1i−N(1+i)N−1

(P/A,g,i,N) = (PA,io,N)⁄1+g

io=1+i1+g−1
Answer #1

A. PV = $ 1000(P/F, 10%,10)

Ру $ 1. 000 *

РУ-$1 . 000 * 0.3855

PV-$385.50

B. The present value of cash flow can be calculated using the following formula

PV = $ 100(P/A,10%,30) + $50(A/G,10%,30)

PV = $ 100* 9.427 + $ 50* 8.176

PV = $ 942.70 + $ 408.80

PV = $ 1351.50

C. The present value of cash flow can be written as follows

A(1 g)

PV_A

Here i=g

Therefore,

large PV = A*rac{n}{1+i}

PV=500 *1+0.1

Py-$9090.91

D. The cash flow can be written as follows

FV = A+ A(1+i)+A(1+i)2+...+A(1+i)44

45 A(1 i)-A FT

FV=A+ (1 + i)45-1

(10.1)45- 1 0.1 1,000,000-A*

large 1,000,000 =A*718.9048

A = $ 1391.005

Every year he must deposit $ 1391.

Please contact if having any query thank you.